# GraspDiffusion: Synthesizing Realistic Whole-body Hand-Object Interaction

### v1

@@ Abstract
Recent generative models can synthesize high-quality images but often fail to generate humans interacting with objects using their hands. This arises mostly from the model’s misunderstanding of such interactions, and the hardships of synthesizing intricate regions of the body. In this paper, we propose GraspDiffusion, a novel generative method that creates realistic scenes of human-object interaction. Given a 3D object mesh, GraspDiffusion first constructs life-like whole-body poses with control over the object’s location relative to the human body. This is achieved by separately leveraging the generative priors for 3D body and hand poses, optimizing them into a joint grasping pose. The resulting pose guides the image synthesis to correctly reflect the intended interaction, allowing the creation of realistic and diverse human-object interaction scenes. We demonstrate that GraspDiffusion can successfully tackle the relatively uninvestigated problem of generating full-bodied human-object interactions while outperforming previous methods. Code and models will be available at https://webtoon.github.io/GraspDiffusion
{strip}[Uncaptioned image]
@@ 1 Introduction
The recent advent of diffusion-based generative models [62, 63, 23] has demonstrated significant success in producing high-quality visual content [58, 51, 57, 61, 54]. When trained on large datasets, these models can coherently synthesize images of various subjects, corresponding to given textual/visual cues. However, despite their strong performance, generative models struggle to comprehend and visualize everyday hand-object interactions. This limitation hinders their broader adoption in generative model-based content creation pipelines.
@@
Located in the terminal regions of the human body, hands occupy only marginal areas within the image distribution, yet have a complex anatomical structure that presents a wide variety of possible hand poses, making them a hard target for image synthesis. Additionally, hands often interact with other objects that also come in various shapes, sizes, and orientations, further complicating their representation. This makes the conditional distribution of hand-object interaction highly complex and convoluted, making it difficult to be solved solely by data acquisition. Consequently, generated images not only suffer from poor hand quality (distorted hand poses, incorrect number of fingers, and uncanny hand shapes) but also exhibit unrealistic interactions (multiple arms from a shoulder, more than one object being interacted, multiple humans being portrayed, etc). Examples of such inaccurate generation are displayed in Fig. 2.
Previous approaches have attempted to handle the issue of correcting hand-generation problems. Several papers [46, 69, 53] have utilized a ControlNet [78] based inpainting approach to refine the appearance of existing hands, but they only focus on situations where the hand is not interacting nor occluded by other objects, making it impractical in many use cases. Other methods [76] directly challenge the synthesis of hand-object interaction by generating a hand for a given object. However, the results are mostly limited in camera diversity and human identity, as the pipelines focus on creating object-centric views displaying a single hand, devoid of a human identity and other implicit interactions (i.e. human gaze towards an object, general body direction).
In this paper, we propose GraspDiffusion, a novel method for generating realistic, whole-bodied interaction images. To ensure the accurate depiction of an interaction, we design a two-stage framework that focuses on utilizing a 3D prior for content creation. In the first stage, we introduce a 3D-context diffusion pipeline that generates the joint human body-object pose in 3D with correct hand-grasping poses, functioning as a concrete scene context. From a 3D object mesh and its position respective to the human body, we separately apply a hand-grasping network [65] and a body-pose diffusion network, combined to produce a 3D human body model naturally interacting with the object. Compared to previous approaches [64, 73, 16, 66, 81], our method successfully generates realistic grasping poses for objects far from the human body, and does not require the modeling of any temporal aspects or extensive test-time optimization, making it effective for practical usage.
In the second stage, we extract guidance from the generated 3D pose to synthesize a realistic image. We train conditional models [78, 49] that accept spatial information to guide the generated image’s structure and appearance. We also apply a cross-attention modulation scheme [3, 13] to control the generation process and prevent unintended interaction. Finally, for additional quality control, we use a conditional inpainting module that can rectify the possibly malformed hand-object interaction, ensuring the perceptual quality is maintained without harming the interaction.
Our GraspDiffusion is the first approach to address the problem of realistic full-bodied interaction synthesis. We demonstrate in our experiments that our method outperforms similar approaches in synthesizing realistic interactions for a given object, displaying plausible hand-object interactions that are both explicit (direct hand grasp) and implicit (torso direction, eye gaze) while preserving the given object identity. The lightweight nature of our pipeline, requiring only the object mesh and its relative position makes our method a viable solution for practitioners in adopting AI-based pipelines.
Figure 2: Negative Examples. We display examples of faulty interaction that were generated using SDXL [54] and ControlNet [78]. Such examples not only display unnatural hands, but also show bizzare object shapes and numbers (e.g. cane separated in half, an axe malformed to fit the human posture, a man holding two coffee cups, a boy reading two books)
@@
Figure 3: GraspDiffusion architecture. We present a two-stage pipeline to generate realistic human-object-interaction images. The first stage takes a single object model and its human-centric location to synthesize a 3D full-bodied grasping pose, providing scene-level context for image generation. The second stage takes reference from the 3D grasping pose, conditionally generating high-quality images.
@@ 2 Related Work
Several papers [46, 69, 53] proposed refining the hands of generated images through Controlnet-based inpainting, but does not count as a direct solution to human-object interaction. Others focused on identifying contacts and inpainting a new hand or object for a given scene [76, 34, 48, 74]. Such solutions, however, are rather limited in camera views and require a prior scene context, making it infeasible in practical scenarios. HOIDiffusion [79] also generates images from synthesized grasps, but is limited to hand-centric views. Compared to HanDiffuser [50], which applied the injection of hand embeddings during image generation to create realistic hands, our pipeline focuses more on the joint synthesis of hand and object, and uses spatial guidance to better direct the generation process towards a 3D scene context.
Grasp Synthesis. Synthesizing a hand grasp consisting of a given object and a hand model is important in understanding human-object interaction, and is a widely studied task in robotics, graphics, and computer vision. While the focus in robotics is to make stable grasps for a given object in simulation / real life [39, 68, 70, 15, 71], in computer vision and graphics the focus is to make plausible grasps that are physically sound, and either generate grasps for hands [12, 25, 28, 82, 80, 11, 43, 40] or a full-bodied grasp [66, 16, 64, 73, 7, 81, 47].
Thanks to the recent advent of human-object interaction datasets [4, 6, 10, 14, 19, 20, 35, 44, 65], many grasp synthesis methods achieved high performance in generating plausible grasps. Most existing datasets, however, have issues with data scalability and variability, hindering the usage of existing datasets in enhancing image generation. While BEHAVE [4] contains both RGB video sequences with 3D annotation, the overall image quality and motion sensors worn by the subject make it hard to use as a realistic image dataset. To overcome this issue, we leverage the traditional human-object interaction datasets [18, 17, 9, 41, 26] along with annotation tools [33, 42, 38, 36, 29] to create pseudo 3D-annotations for the 2D image.
Building our insight from similar approaches, [66, 81], we leverage the priors from a full-body pose model and a hand-grasping model for grasping pose creation. Compared to prior approaches [64, 16, 73, 7], instead of generating a motion sequence, we synthesize the pose parameters for the 3D parametric models [52, 59] using a diffusion model.
@@ 3 GraspDiffusion
Fig. 3 illustrates the proposed architecture. Starting with a 3D object mesh and its position within the human-centric coordinate system (originating at the pelvis joint), GraspDiffusion synthesizes realistic images portraying a human interacting with the object, with a significant portion of the human body visible. We first generate the pose parameters for the human body model grabbing the 3D object mesh (Section. 3.2), from which we extract geometric structures to guide the generation of realistic images, leveraging Stable Diffusion [58] along with spatial encoders and a cross-attention modulation scheme (Section. 3.3).
3.1 Preliminaries.
Diffusion Models. Diffusion models [62, 23] are a group of generative models that interpret the data distribution $p(x)$ as a sequential transformation from a tractable prior distribution $p(x_{T})\sim\mathcal{N}(0,I)$. During training, the model uses a forward noise process $q(x_{t}|x_{t-1})$ that gradually adds a small amount of noise to a clean data sample $x_{0}$ towards $p(x_{T})$. At the same time, the model learns a backward noise process $p(x_{t-1}|x_{t})$ implemented as a neural network, which is trained to remove the noise from the before generating samples from $p(x_{T})$. For Stable Diffusion models [58], the diffusion process is performed in the latent space of a trained autoencoder model [32], guided by a conditional text embedding derived from the CLIP [56] mechanism.
@@
Figure 4: Full Body Grasping Pipeline. The hand-grasp diffusion model takes the BPS [55] representation of the object and generates the MANO [59] parameters for a plausible hand-grasp pose. The body-pose diffusion model takes the object position and generates the SMPL-X [52] parameters for a body-grasp pose. A joint optimization between the poses are made to adjust the object position to a full-bodied grasp pose.
3.2 Full-Body Grasping Pipeline.
Building on prior approaches [66, 81], we separately generate hand grasps and body poses in creating a whole-body grasping pose. Specifically, we take a 3D object mesh, its relative location to the human root, and the contacting hand orientation (left or right) as the input, to generate a SMPL-X mesh that grasps the given 3D object with a realistic body pose and hand-object contact.
@@
Figure 5: Scene Generation Pipeline We first use the rendered object image, skeleton map, joint depth images and the segmentation images as guidance for the conditional generation of a high-quality HOI image. We then use the same types of renderings centered towards the hand-object region for the conditional inpainting to refine the hand quality.
3.3 Scene Generation Pipeline.
Given the 3D whole-body grasping pose, we first extract multiple spatial conditions and leverage the recent Stable Diffusion [58] models to create consistent images of human-object interaction. To further adjust interaction and correct erroneous details (e.g., wrong number of fingers, unnatural textures) we refine the image focused on the hand-object region.
Figure 6: Attention Injection Scheme The hand-grasp diffusion model takes the BPS [55] representation of the object and generates the MANO [59] parameters for a plausible hand-grasp pose. The body-pose diffusion model takes the object position and generates the SMPL-X [52] parameters for a body-grasp pose. A joint optimization between the poses are made to adjust the object position to a full-bodied grasp pose.
To precisely control the human-object image’s generation, we render three spatial conditions from the full-body grasping output. We first render the skeleton $(s^{i})$ of the SMPL-X body consisting of body and hand joints, to ensure realistic human proportions within the generated image. We also use the joint depth map $(d^{i})$ from the SMPL-X and object model to provide depth information. Lastly, we render the occluded object with ambient lighting $(o^{i})$ to preserve its appearance while naturally relighting it. To apply conditions, we chose the CoAdapter [49] approach, which allows flexibility in handling multiple conditions. For each condition, we separately apply an adapter $\mathcal{F}_{\text{AD}}$ and perform a weighted sum to create feature $\mathbf{F}_{c}$, which can be written as
$\mathbf{F}_{c}=\sum_{k\in\{s,d,o\}}\omega_{k}\mathcal{F}_{\text{AD}}^{k}(k^{i}).$ (3)
During training, we fix the parameters in SD and only optimize the conditional adapters, reducing the risk of the model converging to the dataset’s style. Following [49], the training loss becomes:
@@
For hand-object refinement, we utilize the full-body grasping output to produce a joint hand-object mask and spatial conditions: hand skeleton information $(s_{h}^{i})$, a joint hand-object depth map $(d_{h}^{i})$, and the occluded rendered object $(o_{h}^{i})$. These masks and conditions serve as inputs to the conditional inpainting pipeline, refining the structure and appearance of both hand and object while preserving visual integrity. A full illustration of the procedure is in Fig. 5.
@@
4 Experiments.
4.1 Datasets.
To compensate for the lack of realistic, 3D-annotated human-object interaction datasets, we designed an annotation pipeline through which we leveraged previous interaction datasets [18, 17, 9, 41, 26] to function as a pseudo-3D interaction dataset. Specifically, we use the human-object interaction images from HICO-DET [9] and V-COCO [18], which contain a large variety of possible interactions and annotations for the human body and object type.
For both datasets, we first filtered the images so that each image included at least one visible human with a reasonable screen size, along with at least one identifiable hand. To correctly identify the object being interacted with, we use the BLIP-2 language model [37] to perform a Visual Question Answering task, which outputs an object type from the input image. Using the object text, we use GroundingDINO [42] to get the object location and detect the segmentation map of the object, and employ a state-of-the-art depth estimation model [29] to estimate the depth map of the object. Note that for V-COCO, we use the original annotated information for objects. Meanwhile, we also estimate the 3D SMPL-X parameters for the human, using the annotated bounding box and HybrIK [38, 36] 3D human pose and shape estimator. For the human skeleton, we use the annotated Halpe dataset [26] for the HICO-DET dataset and the DWPose estimator [8] for the V-COCO dataset. Among the processed images, we identify images with sufficiently large hands portrayed and reserve them for the hand-object refinement model training. To estimate the 3D MANO parameters, we use the ACR [77] hand pose and shape estimator.
To further augment the dataset, we use the BEHAVE interaction dataset [4] that comes with SMPL-X parameters and object 3D models. In total, we collected 25K joint interaction pairs (image, text, depth map, skeleton, segmentation), which are used to train the scene generation pipeline.
@@
(c) Human skeleton + Object image
@@
Figure 7: Processed HICO-DET [9] dataset sample

### v2

@@ Abstract
Recent generative models can synthesize high-quality images but often fail to generate humans interacting with objects using their hands. This arises mostly from the model’s misunderstanding of such interactions, and the hardships of synthesizing intricate regions of the body. In this paper, we propose GraspDiffusion, a novel generative method that creates realistic scenes of human-object interaction. Given a 3D object mesh, GraspDiffusion first constructs life-like whole-body poses with control over the object’s location relative to the human body. This is achieved by separately leveraging the generative priors for 3D body and hand poses, optimizing them into a joint grasping pose. The resulting pose guides the image synthesis to correctly reflect the intended interaction, allowing the creation of realistic and diverse human-object interaction scenes. We demonstrate that GraspDiffusion can successfully tackle the relatively uninvestigated problem of generating full-bodied human-object interactions while outperforming previous methods.
@@ 1 Introduction
The recent advent of diffusion-based generative models [66, 67, 27] has demonstrated significant success in producing high-quality visual content [62, 55, 61, 65, 58]. When trained on large datasets, these models can coherently synthesize images of various subjects corresponding to given textual/visual cues. However, despite their strong performance, generative models struggle to comprehend and visualize everyday hand-object interactions. This limitation hinders their broader adoption in generative model-based content creation pipelines.
@@
One challenge in generating images with high-quality hands arises from the fact that hands occupy only marginal areas within a full-bodied human image, yet have a complex anatomical structure that presents a wide variety of possible hand poses. Hands come in varying shapes, sizes, orientations, and multiple finger joints that can bend in various degree to support diverse hand poses. Moreover, hands are usually interacting with various objects, making the distribution of hands highly complex and convoluted, leading to faulty generation results (e.g., distorted hand poses, multiple arms from a shoulder, uncanny hand shapes). Examples of such inaccurate generation are displayed in Fig. 2.
While several papers [50, 73, 57] have applied inpainting for hand region refinement, they only focus on situations where the hand is not interacting with other objects, making it impractical for most use cases. In order to generate realistic grasping hands for a given object, the models must understand the semantics and functionalities provided by the object–a concept well known as affordance [18]. While Affordance Diffusion [82] and HOIDiffusion [85] creates an image of a single grasping hand using affordance, these methods only represent the explicit physical contact between a human hand and the object (devoid of any human identity), and fail to convey the spatial / orientational non-contact relationships, making them unsuitable for understanding human affordance. For instance, when using a cell phone, the relationships between the human’s face and torso with respect to the phone should be considered as part of interaction along with the hand touching the phone. This requirement necessitates the development of an pipeline that creates identifiable human-object images where the human’s body is visible, such that the implicit, non-contact relationships are well captured.
In this paper, we present GraspDiffusion, a novel method for generating interaction images with realistic hands and a clear human identity from a single 3D object input. Instead of relying on textual prompts, which tend to be ambiguous in describing complex interactions [54], we first utilize a diffusion model to generate full-bodied grasping poses [56, 63] conditioned on the object and its position. The generated 3D pose parameters are provided as conditions for the next stage to create high-quality image samples. Specifically, we train multiple conditional models [84, 53] and a novel cross-attention modulation scheme [3, 14] to correctly convey the interaction context without harming the diversity of the generated samples. To overcome the lack of a large-scale image-3D pose paired dataset for human object interaction, we also propose a dataset annotation scheme to gather 3D annotations from previous image-based interaction datasets [20, 19, 10, 45, 30].
As far as our knowledge, our pipeline is the first approach to generate full-bodied HOI images from a given object information, such that both explicit and implicit interactions are portrayed in a physically plausible manner. Our experiments on several metrics show that GraspDiffusion outperforms similar approaches in generating high-quality images with realistic interactions and valid 3D grasping poses. We also display cases where GraspDiffusion can be conditioned with a single object image and support various artistic styles, proving its efficacy as a practical solution for AI practitioners in content creation. The contributions of this work are summarized as follows.
  • We propose GraspDiffusion, a novel generative pipeline that synthesizes realistic, full-bodied HOI images.
  • We divide the pipeline into two stages to facilitate both physically plausible body poses along with identity and style diversity. The first stage provides a rich 3D prior to create lifelike interaction poses, which are then provided as 3D conditions for the image generation stage which prodces images with high quality and diversity.
  • We devise an annotation pipeline for an image-3D paired grasping dataset to facilitate the training of HOI image generation.
  • Our experiments in generating realistic images and their paired 3D grasping poses demonstrate the efficacy and substantial performance improvements over previous SOTA methods, providing advanced physical pose plausibility and perceptive visual quality.
(a) ControlNet
(b) HandRefiner
(c) Ours
Figure 2: Comparison between our method and previous approaches on generating HOI images. While previous methods can generate images conditioned on human pose and refine hand shapes, they are prone to erroneous object creation (top row) or faulty interaction synthesis (bottom row).
@@
Figure 3: We present a two-stage pipeline to generate realistic human-object-interaction images. The first stage takes a single object model and its human-centric location to synthesize a 3D full-bodied grasping pose, providing scene-level context for image generation. The second stage takes reference from the 3D grasping pose, conditionally generating high-quality images.
@@ 2 Related Work
Several papers [50, 73, 57] proposed to refine the hands of images generated using Controlnet-based inpainting, but does not count as a direct solution to human-object interaction. Others focused on identifying contacts for inpainting a new hand or object for a given scene [82, 38, 52, 79], yet they are limited in camera views and require a prior scene context, making them infeasible in practical scenarios. HOIDiffusion [85] also generates images from synthesized grasps, but is limited to hand-centric views. Compared to HanDiffuser [54], which applied the injection of hand embeddings during image generation to create realistic hands, our pipeline focuses more on the joint synthesis of hand and object, and uses spatial guidance to better direct the generation process towards a 3D scene context.
Grasp Synthesis. Synthesizing a hand grasp consisting of a given object and a hand model is important in understanding human-object interaction, and is a widely studied task in robotics, graphics, and computer vision. While the focus in robotics is to make stable grasps for a given object in simulation / real life [43, 72, 74, 16, 76], in computer vision and graphics the focus is to make plausible grasps that are physically plausible, and generate grasps for either hands [12, 29, 32, 88, 86, 13, 47, 44] or a human [70, 17, 69, 78, 8, 87, 51].
Thanks to the advent of human-object interaction datasets [4, 7, 11, 15, 21, 23, 39, 48, 68], many grasp synthesis methods achieved high performance in generating plausible grasps. Yet most existing datasets have issues with data scalability and variability, especially when it comes to color image-3D paired datasets. While hand-object interaction datasets like DexYCB [11], ARCTIC [15] and HOI4D [48] consist of richly annotated image data, they are rather focused only on the hand and object recorded in an egocentric manner, devoid of any human identity and spatial / orientational non-contact relationships. Although BEHAVE [4] contains both RGB video sequences with 3D annotation, the image quality and motion sensors worn by the subject make it difficult to use as a realistic image dataset. To overcome this issue, we used traditional human-object interaction datasets [20, 19, 10, 45, 30] along with annotation tools [37, 46, 40, 41, 33] to create pseudo-3D annotations for the 2D image.
Building our insight from similar approaches, [70, 87], we focus on leveraging priors from a full-body pose model and a hand-grasping model. Compared to previous approaches [69, 17, 78, 8], instead of generating a motion sequence, we synthesize the pose parameters for the 3D parametric models [56, 63] using a diffusion model.
@@ 3 GraspDiffusion
Fig. 3 illustrates the proposed architecture. Starting with a 3D object mesh and its position within the human-centric coordinate system (originating at the pelvis joint), GraspDiffusion synthesizes realistic images portraying a human interacting with the object, with a significant portion of the human body visible to be considered ”full-body”. In the initial full-body grasping stage, we generate the pose parameters for the human body model [56] interacting with the 3D object mesh (Section. 3.2). In the scene generation stage, we extract geometric structures from the pose parameters to guide the generation of realistic images, leveraging a latent diffusion model [62] along with spatial encoders and a cross-attention modulation scheme (Section. 3.3).
3.1 Preliminaries
Diffusion Models. Diffusion models [66, 27] are a group of generative models that interpret the data distribution $p(x)$ as a sequential transformation from a tractable prior distribution $p(x_{T})\sim\mathcal{N}(0,I)$. During training, the model uses a forward noise process $q(x_{t}|x_{t-1})$ that gradually adds a small amount of noise to a clean data sample $x_{0}$ towards $p(x_{T})$. At the same time, the model learns a backward noise process $p(x_{t-1}|x_{t})$ implemented as a neural network, which is trained to remove the noise from the before generating samples from $p(x_{T})$. For latent diffusion models [62], the diffusion process is performed in the latent space of a trained autoencoder model [36], guided by a conditional text embedding derived from the CLIP [60] mechanism.
@@ 3.1 Preliminaries
Figure 4: Full-body grasping pipeline. We separately leverage a hand-grasping model [68] and a body-pose diffusion model, and perform a joint optimization into a full-bodied grasping pose.
3.2 Full-Body Grasping Pipeline
Building on prior approaches [70, 87], we separately generate hand grasps and body poses in creating a whole-body grasping pose. Specifically, we take a 3D object mesh, its relative location to the human root, and the contacting hand orientation (left or right) as the input, to generate an SMPL-X mesh that grasps the given 3D object with a realistic body pose and hand-object contact.
@@ 3.2 Full-Body Grasping Pipeline
Figure 5: Scene generation stage. We inject three image conditions and semantic segmentation images as guidance for the generation of a high-quality HOI image. We then use the same types of renderings centered on the hand-object region to refine the hand quality.
3.3 Scene Generation Pipeline
Given the 3D body pose pose, the scene generation pipeline extracts multiple spatial conditions as conditions to a pre-trained [62] model to create consistent images of human-object interaction. Optionally, the pipeline can refine the image focused on the hand-object region to further adjust interaction and correct erroneous details.
Figure 6: Attention Injection Scheme. During inference, we inject the human/object semantic maps into the cross-attention layers as guidance, encouraging the generation process to be focused on the segmented regions. We also apply a negative semantic map for the object to avoid undesired cases where the opposite hand interacts with the object.
To precisely control the human-object image’s generation, we render three spatial conditions from the full-body grasping output. We first render the skeleton $(s^{i})$ of the SMPL-X body, consisting of body and hand joints, to ensure realistic human proportions within the generated image. We also use the joint depth map $(d^{i})$ from the SMPL-X and object model to provide depth information. Lastly, we render the occluded object with ambient lighting $(o^{i})$ to preserve its appearance while relighting it. To apply conditions, we chose the CoAdapter [53] approach, which allows flexibility in handling multiple conditions. For each condition, we separately apply an adapter $\mathcal{F}_{\text{AD}}$ and perform a weighted sum to create feature $\mathbf{F}_{c}$, which can be written as
$\mathbf{F}_{c}=\sum_{k\in\{s,d,o\}}\omega_{k}\mathcal{F}_{\text{AD}}^{k}(k^{i}).\vspace{-0.2cm}$ (3)
During training of the conditioning pipeline, we fix the parameters of the baseline U-Net model and only optimize the conditional adapters, reducing the risk of the model converging to the dataset’s style. This decision allows us to control the image style by applying LoRAs [28] or using finetuned Stable Diffusion models during inference time, making it suitable for real-world application tasks requiring personalized image generation.
@@ 3.3 Scene Generation Pipeline
For hand-object refinement, we utilize the full-body grasping output to produce a joint hand-object mask and spatial conditions: hand skeleton information $(s_{h}^{i})$, a joint hand-object depth map $(d_{h}^{i})$, and the occluded rendered object $(o_{h}^{i})$. These masks and conditions serve as inputs to the hand refinement adapters, which is akin in structure to the body conditioning adapters yet trained separately. This refines the structure and appearance of both hand and object while preserving visual integrity. A full illustration of the procedure is shown in Fig. 5.
@@
4 Experiments
@@
(c) Human skeleton + Object
@@
Figure 7: Processed HICO-DET [10] dataset sample. We jointly extract the joint depth map for the object and human 3D pose, the human skeleton joint map, and semantic segmentation maps.
4.1 Dataset Construction
To compensate for the lack of realistic, 3D-annotated human-object interaction datasets, we designed an annotation pipeline through which we leveraged previous interaction datasets [20, 19, 10, 45, 30] to construct a pseudo-3D interaction dataset. Specifically, we utilized the human-object interaction images from HICO-DET [10] and V-COCO [20], which contain a large variety of possible interactions and annotations for the human body and object type.
For both datasets, we first filtered the images so that each image included at least one visible human with a reasonable screen size, along with at least one identifiable hand. For datasets like HICO-DET, multiple humans within an image might be interacting with multiple objects, making it necessary to correctly identify the main object of interaction. We use the BLIP-2 language model [42] to perform a Visual Question Answering task, which outputs an object type from the input image. Using the object text, we use GroundingDINO [46] to get the object location and detect the segmentation map of the object, and employ a state-of-the-art depth estimation model [33] to estimate the depth map of the object. Note that for V-COCO, we use the original annotated information for objects Meanwhile, we also estimate the 3D SMPL-X parameters for the human, using the annotated bounding box and HybrIK [40, 41] 3D human pose and shape estimator. For the human skeleton, we use the annotated Halpe dataset [30] for the HICO-DET dataset and the DWPose estimator [9] for the V-COCO dataset. Among the processed images, we identify images with sufficiently large hands portrayed and reserve them for the hand-object refinement model training. To estimate the 3D MANO parameters, we use the ACR [83] hand pose and shape estimator. To further augment the dataset, we use the BEHAVE interaction dataset [4] that comes with SMPL-X parameters and object 3D models, and processed it in a similar manner to create joint image interaction pairs (image, text, depth map, skeleton, segmentation).
For the hand refinement modules, we process the Dex-YCB dataset [11], the RHD dataset [90] that also comes with 3D annotations and video data. By processing the MANO parameters [63] and the object 3D models, we also acquired image interaction pairs specifically focused on the hand region. To broaden the hand-object interaction distribution, we include a subset of the preprocessed HICO-DET dataset, cropped on the hand-object bounding box, as additional training data for hand refinement. In total, we collected 25K joint interaction pairs of the full human-object interaction scenes to train the scene generation pipeline, and 15K joint interaction pairs of the hand-object interaction scenes to train the hand refinement pipeline.
Table 1: Quantitative comparison on full-bodied generation